The monthly market basket for consumers consists of pizza, t-shirts, and rent The table below shows market basket quantities and prices for the base year (Year 1) and in the following year Base Year (Year 1) Quantity 20 Price in the Base Year Price in Year 2 Product Pizza T-Shirts Rent $3.50 $25.00 300.00 $4.38 $22.50 $330 00 n Year 1, the CPI for the economy is 5 (Round both answers to one decimal place) In Year 2, ne CPl for this economy is
Answer #1

In year 1,The CPI for the economy is 100.

As year 1 is the base year, the CPI is always 100 is the base year.

In year 2.

Expenditure in base year

20*3.5+4*25+1*300= 70+ 100+300 =470

Expenditure in year 2

20*4.38+4*22.5+1*330= 87.6+90+330=507.6

CPI= 507.6/470*100=108

Ans 108